# Can SAEs reveal and mitigate racial biases of LLMs in healthcare?

### v1

@@ Appendix C Steering
For Figure 6 shows the effect of $\alpha$ on perplexity. For steering white race, we intervene on layer $19$ (latent 2894) and $31$ (latent 13191) for 2B and 9B respectively. This is because we could not locate latents in the middle layer of the models that exclusively activated on white/Caucasian as a race. The “white” latents we found activated on any occurence of the term ‘white’, such as ‘white blood cell’.
@@ D.1 Prompts
Shane is a 57-year-old Caucasian man who is post-op day 5 status-post L4-L5 transforaminal lumbar interbody fusion for severe back and leg pain. There were no surgical complications, and he continues to have low back and leg pain though he endorses steady improvement in his symptoms. When you examine Shane, you confirm that he has no new neurological deficits, and does have persistent pain at his lumbar surgical site. Upon checking the medication administration record, you note that he has required acetaminophen and oxycodone every 6 hours for pain control.
@@
Shane is a 57-year-old Caucasian man who is post-op day 5 status-post L4-L5 transforaminal lumbar interbody fusion for severe back and leg pain. There were no surgical complications, and he continues to have low back and leg pain though he endorses steady improvement in his symptoms. When you examine Shane, you confirm that he has no new neurological deficits, and does have persistent pain at his lumbar surgical site. Upon checking the medication administration record, you note that he has required acetaminophen and oxycodone every 6 hours for pain control.
@@ D.2 Race latents
* Latent 14319 - we manually checked this on clinical summaries as well as inspected the max-activations and description on Neuronpedia - the latent activates more broadly on any ethnicity, not just Russian.
@@ D.3 Intervening on multiple layers
Table 11: Fractional logit difference (FLDD)

### v2

@@ Appendix C Steering
Figure 7 shows the effect of $\alpha$ on perplexity. Table 9 shows the $\alpha$ values used for steering. To steer towards “white”, we intervene on layer $19$ (latent 2894) and $31$ (latent 13191) for 2B and 9B respectively. This is because we could not locate latents in the middle layer of the models that exclusively activated on white/Caucasian as a race. The “white” latents we found activated on any occurence of the term “white”, such as “white blood cell”. We are not sure why this is the case.
@@ D.1 Prompts
Shane is a 57-year-old White man who is post-op day 5 status-post L4-L5 transforaminal lumbar interbody fusion for severe back and leg pain. There were no surgical complications, and he continues to have low back and leg pain though he endorses steady improvement in his symptoms. When you examine Shane, you confirm that he has no new neurological deficits, and does have persistent pain at his lumbar surgical site. Upon checking the medication administration record, you note that he has required acetaminophen and oxycodone every 6 hours for pain control.
@@
Shane is a 57-year-old White man who is post-op day 5 status-post L4-L5 transforaminal lumbar interbody fusion for severe back and leg pain. There were no surgical complications, and he continues to have low back and leg pain though he endorses steady improvement in his symptoms. When you examine Shane, you confirm that he has no new neurological deficits, and does have persistent pain at his lumbar surgical site. Upon checking the medication administration record, you note that he has required acetaminophen and oxycodone every 6 hours for pain control.
@@ D.2 Race latents
Regarding latent 14319 (*): we manually checked this on clinical summaries as well as inspected the max-activations and description on Neuronpedia — the latent activates more broadly on any ethnicity, not just Russian.
@@ D.3 Intervening on multiple layers
Table 11: Fractional logit difference (FLDD).
@@
Appendix E Use of Large Language Models
We used the free versions of Claude and ChatGPT to assist with code for generating plots.